OpenStudy (anonymous):

Please Help! I'll give you a medal! a coin is tossed and a standard number cube is rolled.what is the probability that the coin shows head and the number cube shows a even number. A.1/6 B.1 C.1/4 D. 1/2

OpenStudy (kirbykirby):

\(P(\text{head})=1/2\) \(P(\text{even number})=P(\{ 2,4,6\})=3/6=1/2\) \[P(\text{head AND even number})=P(\text{head})\times P(\text{even number}) \\ =\frac{1}{2}\times \frac{1}{2}\]